# OpenRubrics: Towards Scalable Synthetic Rubric Generation for Reward Modeling and LLM Alignment

### v1

@@ 1 Introduction
Reward modeling is central to reinforcement learning from human feedback (RLHF) and is widely used to align large language models (LLMs) with human preferences (Ouyang et al., 2022, Wu et al., 2023, Bhaskar et al., 2025). By assigning a scalar score (Ouyang et al., 2022) or preference label (Chen et al., 2025b) to each response, reward modeling provides the optimization signal during training and steers the policy LLM toward generating helpful and harmless responses.
While lots of efforts have been paid on RL with verifiable reward (RLVR) (Guo et al., 2025a, Yue et al., 2025), many high-value applications of LLMs, such as long-form question answering, general helpfulness, operate in inherently subjective domains where correctness cannot be sufficiently captured by binary signals. To bridge this gap, rubrics-as-rewards (RaR) (Gunjal et al., 2025) have emerged as a new paradigm for reward modeling. Rubrics include structured natural language criteria that decompose quality into interpretable and measurable dimensions, providing a more consistent and transparent evaluation framework than scalar judgments. For policy models, rubrics also enable optimization to be guided by explicit principles.
Despite their great promise, the construction of high-quality rubrics remains an open challenge. Existing benchmarks (Arora et al., 2025) curate rubrics with the effort from domain experts, which is costly and difficult to scale. Recent works (Huang et al., 2025, Viswanathan et al., 2025, Gunjal et al., 2025) typically generate rubrics via direct prompting LLMs, but those approaches suffer from limited quality control over rubrics and can be prohibitively expensive when relying on commercial APIs.
@@
Our contributions are threefold:
  • Data Contribution: We introduce OpenRubrics, a large-scale and domain-diverse collection of rubrics. This dataset enables both rubric generation models and rubric-informed reward modeling at scale.
  • Methodological Contribution: We distinguish between two fundamental types of rubrics and propose a novel contrastive rubric generation strategy that trains models to produce comprehensive and discriminative rubrics from prompts and response pairs. In addition, we introduce preference–label consistency that improves the quality and reliability of the rubric.
  • Empirical Contribution: We conduct extensive experiments on eight benchmark datasets. Our proposed rubric reward model, Rubric-RM, consistently outperforms strong baselines by 6.8%. Moreover, when integrated into policy optimization, Rubric-RM enables the policy model to achieve strong performance on challenging instruction following and medical benchmarks with an average gain of 2.9%. Case studies further verify the benefits of combining hard rules and principles and show that rubrics help reduce false positives from overly long outputs.
2 Related Work
@@
Standard reward models assign scalar scores to responses by applying a ranking loss between preferred and rejected outputs under the Bradley–Terry framework (Bradley and Terry, 1952, Liu et al., 2025a). To enhance reasoning capability, generative reward models (GenRMs) incorporate synthesized Chains of Thought (CoT), enabling more accurate reward estimation (Ankner et al., 2024, Yu et al., 2025, Zhang et al., 2025b, Shen et al., 2024, Liang et al., 2025). Beyond the pointwise setting, pairwise reward models have been proposed to directly compare multiple responses (Xu et al., 2025, Liu et al., 2025b). More recently, reinforcement learning has been leveraged to further optimize reward models, enabling them to reason explicitly over comparisons and thereby achieve stronger alignment performance (Chen et al., 2025a, b, Whitehouse et al., 2025, Guo et al., 2025b). Orthogonal to these efforts, our work focuses on improving reward modeling quality through the use of structured rubrics. By introducing rubric-based evaluation signals, we complement existing approaches with an additional layer of interpretability that yield performance gains.
@@
Recent work has explored rubrics for both evaluation and alignment. Rubrics provide structured assessments of model generations (Arora et al., 2025, Hashemi et al., 2024, Pathak et al., 2025, Lin et al., 2025), guide instruction following and domain adaptation (Viswanathan et al., 2025, Gunjal et al., 2025), improve safety via rule-based rewards (Mu et al., 2024), and have been combined with verifiable rewards for reasoning tasks (Huang et al., 2025, Zhou et al., 2025). Yet most existing approaches rely on prompting frontier LLMs to generate rubrics, which limits scalability and consistency. Our work introduces a more scalable framework for high quality synthetic rubric generation, improving both reward quality and interpretability at a cheaper cost. Concurrently, Zhang et al. (2025a) also investigate rubric generation, but focus on iterative refinement to mitigate reward over optimization, whereas we emphasize scalable synthesis and rubric–preference consistency.
@@ 3 Preliminaries
Rubrics.
We define rubrics as a structured set of evaluation criteria tailored to a given prompt. Formally, let $x$ denote an input prompt and $\hat{y}$ a model-generated response. A rubric $\mathcal{R}(x)$ is represented as a collection of $k$ criteria $\mathcal{R}(x)=\{c_{i}\}_{i=1}^{k}$, where each $c_{i}$ denotes a rubric description specifying an aspect of response quality (e.g., factual correctness, reasoning soundness, style).
Rubrics-based Reward Models.
Building on prior work in pairwise reward modeling (Liu et al., 2025b, Chen et al., 2025b, Guo et al., 2025b, Xu et al., 2025), we focus on a comparative setting where the goal is to evaluate the relative quality of two candidate responses. Given a prompt $x$ and two samples $(\hat{y}_{1},\hat{y}_{2})$, a pairwise rubric-based reward function is defined as
$\text{reward}_{\text{pair}}(x,\hat{y}_{1},\hat{y}_{2})=r_{\theta}\big(x,\hat{y}_{1},\hat{y}_{2};\{c_{i}\}_{i=1}^{k}\big),$ (1)
where reward is the binary preference label, $r_{\theta}$ is the reward model parameterized by $\theta$ that integrates rubric criteria $\{c_{i}\}$ when producing a preference judgment.
The overall framework for OpenRubrics is in Figure 1. Our overall objective is two-fold: (i) constructing a rubric dataset $\mathcal{D}_{\text{rubric}}$ for training a generation model $g_{\theta}$ that automatically synthesizes rubrics $\mathcal{R}(x)$ given a prompt $x$; and (ii) building a reward modeling dataset $\mathcal{D}_{\text{rm}}$ for training a rubric-guided reward model $r_{\phi}$ capable of producing reliable and interpretable pairwise judgments. This two-stage formulation enables explicit decomposition of evaluation into rubric generation and rubric-conditioned reward prediction, bridging human-aligned criteria and automated preference modeling.
@@ 4.1 Data Construction
Data Sources. To generate high-quality rubrics that generalize across tasks and domains, it is crucial to construct a dataset with broad coverage in domains and tasks. To this end, we integrate a range of publicly available preference and instruction-tuning datasets, balancing general-purpose conversational data with domain-specific resources. Specifically, our dataset draws from the following sources:
  • UltraFeedback (Cui et al., 2024), which aggregates preference annotations from Evol-Instruct (Xu et al., 2024), UltraChat (Ding et al., 2023), ShareGPT (Chiang et al., 2023), and TruthfulQA (Lin et al., 2022).
  • Tulu 2.5 (Ivison et al., 2024), a diverse mixture of preference datasets including AlpacaFarm (Dubois et al., 2023), Chatbot Arena (Chiang et al., 2024), Capybara [^1] https://huggingface.co/datasets/argilla/distilabel-capybara-dpo-7k-binarized, StackExchange (Lambert et al., 2023), Nectar (Zhu et al., 2023), SHP (Ethayarajh et al., 2022), HH-RLHF (Bai et al., 2022), HelpSteer (Wang et al., 2024c), and Orca (Mukherjee et al., 2023).
  • HelpSteer3 (Wang et al., 2025), a large-scale human-annotated dataset designed for alignment with helpfulness preferences.
  • Skywork-Preference (Liu et al., 2024), which integrates data from HelpSteer2 (Wang et al., 2024b) and OffsetBias (Park et al., 2024).
  • Tulu3-IF (Lambert et al., 2025a), a collection of human preference judgments tailored for verifiable instruction-following.
  • MegaScience (Fan et al., 2025), a domain-specialized corpus spanning multiple scientific domains including physics, medicine, biology, and chemistry.
  • Medical-o1 (Chen et al., 2024), a medical SFT dataset curated for diagnostic reasoning tasks.
Preference Data Construction. To build preference data for rubric generation and judge training (see Sec. 4.3), we reuse existing preference and SFT datasets with tailored processing. For UltraFeedback, we select the highest-scoring response as the chosen and the lowest as the rejected. For Tulu3-IF, MegaScience, and Medical-o1, we generate multiple responses using Qwen-3-8B/14B (Yang et al., 2025), Llama-3.1-8B (Grattafiori et al., 2024), and Gemma-3-12B (Team et al., 2025), selecting one from each model. For Verifiable-IF, responses satisfying all verification functions are labeled as chosen, and others as rejected. For the MegaScience and Medical-o1 datasets, we employ an ensemble of open-source reward models: Athene-RM-8B (Frick et al., 2024a) and Skywork-Reward-V2-Llama-3.1-8B-40M (Liu et al., 2025a) to rank responses and form best–worst preference pairs.
4.2 Rubrics Synthesis
After collecting a diverse set of preference pairs, our next objective is to construct a set of rubrics that serve as anchors for guiding LLM-based reward modeling. To comprehensively represent different types of constraints while preserving discriminative granularity, we categorize rubrics into two complementary types: (i) Hard Rules, which capture explicit requirements stated in the user’s prompt; and (ii) Principles, which describe higher-level qualitative aspects such as reasoning soundness, factuality, or stylistic coherence. We then introduce two strategies for generating high-quality rubrics, detailed as follows:
Contrastive Rubric Generation.
Given a preference dataset $\mathcal{D}=\left\{\left(x_{i},\hat{y}_{i}^{+},\hat{y}_{i}^{-}\right)\right\}_{i=1}^{N},$ where $x_{i}$ is the prompt and $\hat{y}_{i}^{+}$ and $\hat{y}_{i}^{-}$ denote the preferred and displeased responses, respectively. Our objective is to infer rubrics $\mathcal{R}\left(x_{i}\right)$ that capture qualities a good response should satisfy and the criteria by which one response is preferred over another, using $\hat{y}_{i}^{+}$ and $\hat{y}_{i}^{-}$ as guidance. Formally, we prompt an instruction-tuned LLM $h_{\psi}$ as follows:
$\mathcal{R}\left(x_{i}\right)\sim h_{\psi}\left(x_{i},\hat{y}_{i}^{+},\hat{y}_{i}^{-},\ell_{i}\right),$
where $\ell_{i}$ is the preference label. The generator is asked to produce a set of discriminative evaluation criteria $\mathcal{R}\left(x_{i}\right)=\left\{c_{i,1},\ldots,c_{i,k_{i}}\right\}$, where each $c_{i,j}$ describes a specific aspect. This contrastive setting encourages the model to discover rubric dimensions that are both task-sensitive and preference-aligned.
(a) The data distribution for OpenRubrics.
@@ 4.2 Rubrics Synthesis
Rubric Filtering with Preference-label Consistency.
Not all generated rubrics faithfully capture the human preference signal. To ensure reliability, we conduct a consistency-based filtering step via prompting the LLM $h_{\psi}$ again. For each triplet ($x_{i},\hat{y}_{i}^{+},\hat{y}_{i}^{-}$), we concatenate the full rubric $\mathcal{R}\left(x_{i}\right)$ into the context and ask the model to predict which response better aligns with the rubric $\hat{l}_{i}=h_{\psi}\left(x_{i},\mathcal{R}\left(x_{i}\right),\hat{y}_{i}^{+},\hat{y}_{i}^{-}\right),$ where $\hat{l}_{i}=(\hat{r}_{i},\hat{\ell}_{i})$ is the final prediction, $\hat{r}_{i}$ denotes the prediction rationale and $\hat{\ell}_{i}$ denotes the predicted preference. We retain only those rubrics that lead to predictions consistent with the original human label $\ell_{i}$:
@@
This yields a filtered collection of high-quality rubrics that are both interpretable and empirically consistent with human preferences. The final rubrics-conditioned preference dataset combines prompts, paired responses, and their verified rubrics as:
$\mathcal{D}_{\text{rubric }}=\left\{\left(x_{i},\hat{y}_{i}^{+},\hat{y}_{i}^{-},\mathcal{R}^{*}(x_{i})\right)\right\}_{i=1}^{M}.$
Figure 3: The T-SNE plot for the embeddings of prompts.
Rubric Statistics Overview.
We analyze the curated rubric set along three axes: (i) domain coverage (instruction following, reasoning, general helpfulness; Figure 2(a)); (ii) the balance between hard rules and principles as well as the length of prompts and rubrics (Figure 2(b)); and (iii) semantic diversity of prompt topics, visualized via t-SNE on embeddings from Qwen-3-Embedding-0.6B (Zhang et al., 2025c) (Figure 3).
These statistics confirm that the synthesized rubrics provide comprehensive, yet discriminative coverage, forming a foundation for rubric-based reward modeling.
@@ 4.3 Reward Model Training and Inference
After collecting the rubrics-based dataset, we proceed to develop a rubric generation model that outputs systematic evaluation rubrics and a reward model Rubric-RM that generates final preference labels.
Rubric Generation.
We first train a rubric generation model $g_{\theta}$ to produce structured evaluation rubrics conditioned on the prompt and paired responses. Formally, given a dataset $\mathcal{D}_{\text{rubric}}=\{(x_{i},\hat{y}_{i}^{+},\hat{y}_{i}^{-},\mathcal{R}^{*}(x_{i}))\}_{i=1}^{M}$, where $\mathcal{R}^{*}(x_{i})$ denotes the reference rubric associated with the prompt, the model $g_{\theta}$ is trained via supervised fine-tuning (SFT) with the standard next-token cross-entropy loss:
@@
This objective teaches the model to generate detailed, domain-relevant rubrics that encode evaluation criteria from the user-provided prompt, which can be subsequently used in reward modeling.
Reward Model Training.
Using the synthesized rubrics, we then train the reward model $r_{\phi}$ on $\mathcal{D}_{\text{rm}}=\{(x_{i},\hat{y}_{i}^{+},\hat{y}_{i}^{-},\mathcal{R}^{*}(x_{i}),\hat{l}_{i})\}_{i=1}^{M}$. The model is also optimized via SFT to predict the label tokens $\hat{l}_{i}$ conditioned on the prompt, response pair, and rubric:
@@
Inference.
At inference time, given a pairwise test instance $(x,y^{\text{A}},y^{\text{B}})$, Rubric-RM performs a two-stage process to predict the final preference label: (1) the rubric generator first produces $\hat{\mathcal{R}}(x)=g_{\theta}(x,y^{\text{A}},y^{\text{B}})$; (2) the reward model then predicts the verdict conditioned on the generated rubric from two possible labels $\mathcal{C}=\{\text{A is better},\text{B is better}\}$:
@@
This pipeline ensures that the judgment from Rubric-RM is explicitly grounded in rubric criteria.

### v2

@@ 1 Introduction
Reward modeling is central to reinforcement learning from human feedback (RLHF) and is widely used to align large language models (LLMs) with human preferences (ouyang2022training, wu2023fine, bhaskar2025language). By assigning a scalar score (ouyang2022training) or preference label (chen2025rm) to each response, reward modeling provides the optimization signal during training and steers the policy LLM toward generating helpful and harmless responses.
While lots of efforts have been paid on RL with verifiable reward (RLVR) (guo2025deepseek, yue2025does), many high-value applications of LLMs, such as long-form question answering, general helpfulness, operate in inherently subjective domains where correctness cannot be sufficiently captured by binary signals. To bridge this gap, rubrics-as-rewards (RaR) (gunjal2025rubrics) have emerged as a new paradigm for reward modeling. Rubrics include structured natural language criteria that decompose quality into interpretable and measurable dimensions, providing a more consistent and transparent evaluation framework than scalar judgments. For policy models, rubrics also enable optimization to be guided by explicit principles.
Despite their great promise, the construction of high-quality rubrics remains an open challenge. Existing benchmarks (arora2025healthbench) curate rubrics with the effort from domain experts, which is costly and difficult to scale. Recent works (huang2025reinforcement, viswanathan2025checklists, gunjal2025rubrics) typically generate rubrics via direct prompting LLMs, but those approaches suffer from limited quality control over rubrics and can be prohibitively expensive when relying on commercial APIs.
@@
Our contributions are three-fold:
  • We introduce OpenRubrics, a large-scale and diverse collection of rubrics. This dataset enables both rubric generation models and rubric-informed reward modeling at scale.
  • We distinguish between two fundamental types of rubrics and propose a contrastive rubric generation strategy that trains models to produce comprehensive and discriminative rubrics from prompts and responses. Besides, we introduce preference–label consistency that improves the quality and reliability of the rubric.
  • We conduct extensive experiments on eight benchmark datasets, where Rubric-RM consistently outperforms strong baselines by 8.4%. Moreover, when integrated into policy optimization, Rubric-RM consistently yields notable gains on challenging instruction following and medical benchmark. Case studies further verify the benefits of combining hard rules and principles, showing that rubrics help reduce false positives from overly long outputs.
2 Related Works
@@
Standard reward models assign scalar scores to responses by applying a ranking loss under the Bradley–Terry framework (bradley1952rank, liu2025skywork). To enhance reasoning capability, generative reward models (GenRMs) incorporate synthesized Chains of Thought (CoT), enabling more accurate reward estimation (ankner2024critique, yu-etal-2025-self, zhang2025generative, liang-etal-2025-generative). Beyond the pointwise setting, pairwise reward models have been proposed to directly compare multiple responses (xu2025unified, liu2025pairjudge). More recently, reinforcement learning has been leveraged to further optimize reward models, enabling them to reason explicitly over comparisons and thereby achieve stronger alignment performance (chen2025judgelrm, chen2025rm, whitehouse2025j1, guo2025reward). Orthogonal to these efforts, our work focuses on improving reward modeling quality with structured rubrics. By introducing rubric-based evaluation signals, we complement existing approaches with an additional layer of interpretability that yield performance gains.
@@
Recent work has explored rubrics for both evaluation and alignment. Rubrics provide structured assessments of model generations (arora2025healthbench, hashemi2024llm, pathak2025rubric, lin2025wildbench), guide instruction following and domain adaptation (viswanathan2025checklists, gunjal2025rubrics), improve safety via rule-based rewards (mu2024rule), and have been combined with verifiable rewards for reasoning tasks (huang2025reinforcement, zhou2025breaking). Yet most existing approaches rely on prompting frontier LLMs to generate rubrics, which limits scalability and consistency. Our work introduces a more scalable framework for high quality synthetic rubric generation, improving both reward quality and interpretability at a cheaper cost. Concurrently, zhang2025chasing also investigate rubric generation, but focus on iterative refinement to mitigate reward over optimization, whereas we emphasize scalable synthesis and rubric–preference consistency.
@@ 3 Preliminaries
Rubrics. We define rubrics as a structured set of evaluation criteria tailored to a given prompt. Formally, let $x$ denote an input prompt and $\hat{y}$ a model-generated response. A rubric $\mathcal{R}(x)$ is represented as a collection of $k$ criteria $\mathcal{R}(x)=\{c_{i}\}_{i=1}^{k}$, where each $c_{i}$ denotes a rubric description specifying an aspect of response quality (e.g., factual correctness, reasoning soundness, style).
Rubrics-based Reward Models. Building on prior work in pairwise reward modeling (liu2025pairjudge, chen2025rm, guo2025reward, xu2025unified), we focus on a comparative setting where the goal is to evaluate the relative quality of two candidate responses. Given a prompt $x$ and two samples $(\hat{y}_{1},\hat{y}_{2})$, a pairwise rubric-based reward function is defined as
$\text{reward}_{\text{pair}}(x,\hat{y}_{1},\hat{y}_{2})=r_{\theta}\big(x,\hat{y}_{1},\hat{y}_{2};\{c_{i}\}_{i=1}^{k}\big),$
where reward is the binary preference label, $r_{\theta}$ is the reward model that integrates rubric criteria $\{c_{i}\}$ when producing a preference judgment.
The overall framework for OpenRubrics is in Figure 1. Our overall objective is two-fold: (1) constructing a rubric dataset $\mathcal{D}_{\text{rubric}}$ to train a generation model $g_{\theta}$ that automatically synthesizes rubrics $\mathcal{R}(x)$ given a prompt $x$; and (2) building a reward modeling dataset $\mathcal{D}_{\text{rm}}$ to train a rubric-guided reward model $r_{\phi}$ capable of producing reliable and interpretable pairwise judgments. This two-stage formulation decomposes evaluation into rubric generation and rubric-conditioned reward prediction, bridging human-aligned criteria and automated preference modeling.
@@ 4.1 Data Construction
Data Sources. To generate high-quality rubrics that generalize across tasks and domains, we integrate a range of publicly available preference and instruction-tuning datasets, balancing general-domain data with domain-specific resources. Specifically, our dataset draws from the following sources:
  • UltraFeedback (cui2024ultrafeedback), which aggregates preference annotations from Evol-Instruct (xu2024wizardlm), UltraChat (ultrachat), ShareGPT (sharegpt), and TruthfulQA (lin-etal-2022-truthfulqa).
  • Magpie (xu2025magpie): a large-scale synthetic alignment dataset generated by self-prompting LLMs across diverse domains.
  • Skywork-Preference (liu2024skywork), which integrates data from HelpSteer2 (wang2024helpsteer) and OffsetBias (offsetbias).
  • Synthetic-IF (lambert2025tulu), a collection of human preference judgments tailored for verifiable instruction-following.
  • MegaScience (fan2025megascience), a specialized corpus spanning multiple scientific domains including physics, biology, and chemistry.
  • Medical-o1 (chen2024huatuogpt), a medical SFT dataset curated for diagnostic reasoning tasks.
(a) The data distribution for OpenRubrics (in # instructions).
@@
Preference Data Construction. To build preference data for rubric generation and judge training (see Sec. 4.3), we reuse existing preference and SFT datasets with tailored processing. For UltraFeedback, we select the highest-scoring response as the chosen and the lowest as the rejected. For MegaScience, and Medical-o1, we generate multiple responses using Qwen-3-8B/14B (yang2025qwen3), Llama-3.1-8B (grattafiori2024llama), and Gemma-3-12B (team2025gemma), selecting one from each model. For Synthetic-IF, responses satisfying all verification functions are labeled as chosen, and others as rejected. For the MegaScience and Medical-o1 datasets, we employ an ensemble of open-source reward models: Athene-RM-8B (Athene2024) and Skywork-Reward-V2-Llama-3.1-8B-40M (liu2025skywork) to rank responses and form best–worst preference pairs.
4.2 Rubrics Synthesis
After collecting a diverse set of preference pairs, our objective is to construct a set of rubrics that serve as anchors to guide reward modeling. To comprehensively represent different types of constraints while preserving discriminative granularity, we categorize rubrics into two types: (1) Hard Rules, which capture explicit requirements stated in the user’s prompt; and (2) Principles, which describe higher-level qualitative aspects such as reasoning soundness, factuality, or stylistic coherence. We then introduce two strategies for generating high-quality rubrics as follows:
Contrastive Rubric Generation. Given a preference dataset $\mathcal{D}=\left\{\left(x_{i},\hat{y}_{i}^{+},\hat{y}_{i}^{-}\right)\right\}_{i=1}^{N},$ where $x_{i}$ is the prompt and $\hat{y}_{i}^{+}$ and $\hat{y}_{i}^{-}$ denote the preferred and displeased responses, respectively. Our objective is to infer rubrics $\mathcal{R}\left(x_{i}\right)$ that capture qualities a good response should satisfy and the criteria by which one response is preferred over another, using $\hat{y}_{i}^{+}$ and $\hat{y}_{i}^{-}$ as guidance. Formally, we prompt an instruction-tuned LLM $h_{\psi}$ as $\mathcal{R}\left(x_{i}\right)\sim h_{\psi}\left(x_{i},\hat{y}_{i}^{+},\hat{y}_{i}^{-},\ell_{i}\right),$ where $\ell_{i}$ is the preference label. The generator is asked to produce a set of discriminative evaluation criteria $\mathcal{R}\left(x_{i}\right)=\left\{c_{i,1},\ldots,c_{i,k_{i}}\right\}$, where each $c_{i,j}$ describes a specific aspect. This setting encourages the model to discover rubric dimensions that are both task-sensitive and preference-aligned.
Rubric Filtering with Preference-label Consistency. Not all generated rubrics faithfully capture the human preference signal. To ensure reliability, we conduct a consistency-based filtering step via prompting the LLM $h_{\psi}$ again. For each triplet ($x_{i},\hat{y}_{i}^{+},\hat{y}_{i}^{-}$), we concatenate the full rubric $\mathcal{R}\left(x_{i}\right)$ into the context and ask the model to predict which response better aligns with the rubric $\hat{l}_{i}=h_{\psi}\left(x_{i},\mathcal{R}\left(x_{i}\right),\hat{y}_{i}^{+},\hat{y}_{i}^{-}\right),$ where $\hat{l}_{i}=(\hat{r}_{i},\hat{\ell}_{i})$ denotes the prediction rationale and the predicted preference, respectively. For pairwise judging, we retain only those rubrics that lead to predictions consistent with the original human label $\ell_{i}$:
@@ 4.2 Rubrics Synthesis
For listwise judging settings, where a single prompt is associated with multiple candidate responses, we decompose the task into multiple pairwise comparisons. In this case, we retain a rubric if it yields preference predictions consistent with human labels in at least $50\%$ of the induced pairwise comparisons, allowing robustness to occasional ambiguity while preserving discriminative power.
This yields a collection of high-quality rubrics that are both interpretable and empirically consistent with human preferences. The final rubrics-conditioned preference dataset combines prompts, paired responses, and their verified rubrics as $\mathcal{D}_{\text{rubric }}=\left\{\left(x_{i},\hat{y}_{i}^{+},\hat{y}_{i}^{-},\mathcal{R}^{*}(x_{i})\right)\right\}_{i=1}^{M}.$
Figure 3: The T-SNE plot for embeddings of prompts.
Rubric Statistics Overview. We analyze the curated rubric set along three axes: (1) domain coverage (instruction following, reasoning, general helpfulness; Figure 2(a)); (2) the balance between hard rules and principles as well as the length of prompts and rubrics (Figure 2(b)); and (3) semantic diversity of prompt topics, visualized via t-SNE on embeddings from Qwen-3-Embedding-0.6B (zhang2025qwen3) (Figure 3). These statistics confirm that the synthesized rubrics provide comprehensive, yet discriminative coverage, forming a foundation for rubric-based reward modeling.
@@ 4.3 Reward Model Training and Inference
After collecting the rubrics-based dataset, we proceed to develop a rubric generation model that outputs evaluation rubrics and a reward model Rubric-RM that generates final preference labels.
Rubric Generation. We first fine-tune $g_{\theta}$ to generate rubrics $\mathcal{R}^{*}$ conditioned on the prompt $x$. Given the dataset $\mathcal{D}_{\text{rubric}}=\{(x_{i},\hat{y}_{i}^{+},\hat{y}_{i}^{-},\mathcal{R}^{*}(x_{i}))\}_{i=1}^{M}$, we optimize the standard cross-entropy objective:
@@
Reward Model Training. We then train the reward model $r_{\phi}$ to predict preference labels $\hat{l}$ using the generated rubrics. Using $\mathcal{D}_{\text{rm}}=\{(x_{i},\hat{y}_{i}^{+},\hat{y}_{i}^{-},\mathcal{R}^{*}(x_{i}),\hat{l}_{i})\}_{i=1}^{M}$, we optimize $r_{\phi}$ to predict $\hat{l}$ conditioned on the prompt, response pair, and rubric:
@@
Inference. At inference time, given a pairwise test instance $(x,y^{\text{A}},y^{\text{B}})$, Rubric-RM performs a two-stage process to predict the final preference label: (1) the rubric generator produces (or retrieve a cached) rubric conditioned on the instruction $x$ as $\hat{\mathcal{R}}(x)=g_{\theta}(x)$; (2) the reward model then predicts the verdict over candidate responses $(y^{\text{A}},y^{\text{B}})$ conditioned on the generated rubric from two possible labels $\mathcal{C}=\{\text{A is better},\text{B is better}\}$:
@@
This ensures that the judgment from Rubric-RM is explicitly grounded in rubric criteria.